Hello! A coin is flipped three times. Let A be the event that the outcome of the first flip is a heads. Let B be the be the event that the outcomes of second and third flips are both tails..Show formula Please help me!
Answer by DrBeeee(684)

The P(htt) = (1/2)*(1/2)*(1/2) or
P(htt) = 1/8
In fact the probability for any of the 8 possible sequence of 3 coin flips is 1/8.
That is: P(hhh) = P(hht) = P(hth) = P(htt) = P(thh) = P(tht) = P(tth) = P(ttt) = 1/8
This is because it is equally likely to flip a head or tail.